### v2

Abstract
This paper presents a secure and versatile sampling-based verification protocol, Proof of Sampling (PoSP) protocol, suitable for a wide range of decentralized applications. Our protocol has a unique Nash Equilibrium in pure strategies, which compels rational participants to act honestly, thus fortifying the network’s integrity. This can be achieved with manageable computational overhead. When applied to decentralized inference for AI applications, we design spML based on PoSP protocol, which ingeniously amalgamates the strengths of optimistic fraud proof and zero knowledge proof based approaches, the foremost approaches in the domain at present. Moreover, the PoSP protocol can be effectively utilized for designing verification mechanisms within Actively Validated Services (AVS) in EigenLayer and Layer 2 solutions, further broadening its applicability. This innovative approach not only enhances the security and efficiency of decentralized systems but also paves the way for a new generation of scalable and reliable decentralized applications.
1 Introduction
In the development of decentralized protocols, it is customary to presuppose that honest nodes will adhere to the established protocol. Take, for instance, optimistic rollup [7] is a scaling solution for blockchain that aims to increase blockchain’s transaction throughput. In this approach, a designated rollup validator processes transactions off-chain and posts the results on the blockchain. By default, the transactions are assumed to be executed correctly. However, if other validators detect incorrect transactions, they can submit fraud proofs on-chain to challenge the validator who submitted the false results. The security of optimistic rollup relies on the critical assumption that at least one rollup validator is honest. If all validators are dishonest, the on-chain transactions could be fraudulent. To incentivize honest behavior, an economic reward system is introduced, making it more profitable for rational nodes to act honestly. However, under this assumption, [14] observed that in most existing decentralized systems, the equilibrium state corresponds to a mixed strategy Nash Equilibrium, which means that the optimal strategy for each validator is to cheat a certain percentage of the time! This implies a non-negligible probability of dishonest behavior that is never detected and introduces significant security vulnerabilities. Furthermore, [14] also proposed an alternate framework for optimistic rollups that mandates the scrutiny of the system by all nodes. However, such a design introduces duplicated work from all nodes and contravenes the fundamental principle of scalability, undermining the initial objective of Layer 2 design.
In this paper, we introduce the Proof of Sampling (PoSP) protocol designed to address these challenges effectively, under certain foundational assumptions. The PoSP protocol is applicable across a broad range of decentralized systems. In this paper, we introduce its integration into decentralized AI inference platforms. Notably, our system achieves a unique Nash Equilibrium in pure strategies, wherein each node acts honestly, an principle that is highlighted in [9] as ideal for the design of secure decentralized systems. This signifies that even when each node’s strategy is directed towards maximizing its individual profit, the overall system maintains optimal security.
1.1 Main Contribution
The main contributions of this paper are twofold. First, we propose the PoSP protocol, a robust protocol for a wide range of decentralized applications. PoSP protocol leverages a unique Nash Equilibrium in pure strategies to incentivize rational actors within the network to maintain honest behavior, thereby significantly reinforcing the network’s security and integrity. Second, we design spML, a sampling-based verification mechanism for decentralized AI inference network, using PoSP protocol. SpML synthesizes the advantages of both optimistic fraud proof and zero knowledge proof based approaches (opML and zkML), incorporating the high scalability and simplicity of opML with the enhanced security features of zkML. SpML stands out for its strategic balance of computational efficiency and security, positioning it as a superior alternative within the decentralized AI inference landscape.
In sum, our work not only introduces PoSP as a foundational protocol for secure and efficient verification in decentralized systems but also presents the development of spML, a balanced approach in decentralized AI, derived from the PoSP protocol. These innovations significantly advance the field, fostering a new generation of decentralized applications.
1.2 Design Principles
PoSP enhances decentralized system security by initially selecting an asserter who submits results without knowing which validators might review them. If challenged, $n$ validators are randomly selected, and only then do they learn each other’s identities. This approach prevents collusion at the outset and limits free-riding by concealing the results until all verifications are complete, ensuring independent validation by each participant.
In the case where all nodes agree on the outcome, the system accepts this consensus, ensuring that the protocol functions correctly. However, if there are discrepancies, indicating dishonesty or error, those nodes are penalized. This mechanism ensures that acting honestly is not only the most straightforward but also the most beneficial strategy for all participants, thus maintaining the integrity and security of the system.
1.3 Literature Review
Employing game theory for protocol analysis, rational and economic theories have been extensively applied to explore various blockchain configurations, including Byzantine Fault Tolerance as discussed in [1, 6], sharding strategies [15], proof-of-work systems [2, 5, 10, 13], proof-of-stake systems [3, 16, 17], secure outsourced computation [20], and Layer 2 solutions [11, 12, 14, 18]. The use of slashing mechanisms, analyzed in [9], enhances network security and ideally aims to align node incentives in such a way that following the protocol-prescribed strategy always yields the most benefit, reflecting a unique Nash Equilibrium in pure strategies. A comprehensive review [21] elaborates on the diverse applications of game theory across different blockchain scaling solutions. This analytical approach underscores the utility of game theory in dissecting and enhancing the security and scalability mechanisms of blockchain technology.
The decentralized AI domain has significantly benefited from innovations like opML, as detailed by [3] and [19], which enhances machine learning (ML) on the blockchain by facilitating the efficient processing of intricate models, alongside the introduction of Zero-Knowledge Machine Learning (zkML) by [22, 23]. These advancements in blockchain AI, opML and zkML, address scalability, security, and efficiency with distinct trade-offs. OpML enhances scalability and efficiency, yet its security may not be as robust. ZkML offers strong security through zero-knowledge proofs, yet faces challenges in scalability and efficiency due to high computational overhead.
The rest of this paper is organized as follows. In Section 2, we introduce PoSP protocol and prove that it has a unique Nash Equilibrium in pure strategies. PoSP protocol can be applied in all kinds of verification scenarios. In Section 3, we show a possible way to implement PoSP protocol in real world applications as an example. We design a sampling-based verification mechanism (spML), which is implemented by PoSP protocol, within a decentralized AI inference network, detailing the protocol design and its security validation. In Section 4, we conclude the paper and discuss the future extensions of our work.
2 Model
2.1 PoSP Protocol
In this section, we propose the PoSP protocol.
  1. 1.
    A node is randomly selected from the network to serve as an asserter. This asserter calculates a value $f(x)$, with both the function $f(\cdot)$ and the input $x$ being well-known to the network, and outputs the result.
  2. 2.
    With a predetermined probability $p$, a challenge mechanism protocol is triggered. If the challenge mechanism is not triggered, this round concludes, and the asserter is awarded a reward denoted by $R_{A}$.
  3. 3.
    If the challenge mechanism is triggered, $n$ validators are randomly selected from the network, where $n\geq 1$ is a predetermined integer parameter. Each validator, denoted as validator $i$, independently computes $f(x)$ and outputs the result.
  4. 4.
    If all of the results from the asserter and the valitors match, the result is deemed valid and accepted. The asserter receives $R_{A}$ and each validator receives $R_{V}/n$. Otherwise, an arbitration process is initiated to determine the honesty of the asserter and each of the $n$ validators. If the asserter is proven honest, it receives a reward of $R_{A}$. Otherwise, it is penalized with $S$. Each honest validator, out of the $m$ honest ones from a total of $n$ validators, receives $R_{V}/m$ if $m\geq 1$, while each dishonest validator is penalized with $S$.
2.2 Assumptions and Analysis
Then we show that, under certain conditions, PoSP protocol has a unique Nash Equilibrium in pure strategies that all nodes act honestly. This ensures there is no economic incentive for nodes to perform malicious actions. To analyze this protocol, we additionally define
  • $C$: computational cost for computing $f(x)$
  • $U_{1}$: maximum profit that the asserter can gain if he acts dishonestly and the challenge mechanism is not triggered
  • $U_{2}$: maximum profit that the asserter can gain if the challenge mechanism is triggered and he colludes with all validators
Assumption 1
We assume that if the validator does not collude with the asserter, their messages will never match with each other unless they both act honestly in the network.
Assumption 2
We also assume that $R_{A}-C>-S$ and $R_{V}/n-C>-S$, which means that it is better for each node to receive the reward than be slashed.
Property 1
Under Assumption 1 and Assumption 2, if one validator does not collude with the asserter, his dominant strategy is to act honestly.
Strategy Asserter Fraud Asserter Not Fraud
Validator Honest $R_{V}/m-C$ $R_{V}/m-C$
Validator Dishonest $-S$ $-S$
The table above gives a game in a bimatrix format. The first row is the behavior of the asserter, and the first column is the behavior of the validator, who does not collude with the asserter. The value in the table is the payoff of the validator. In this case, Property 1 is straightforward by the table.
Strategy Detected Undetected
Assserter Fraud $-S$ $U_{1}$ or $U_{2}$
Asserter Not Fraud $R_{A}-C$ $R_{A}-C$
The table above shows the payoff of the asserter. The first row represents whether his behavior will be detected by the protocol, and the first column denotes the behavior of the asserter.
Assumption 3
We assume that the fraction of the nodes in the network that the asserter controls is $r$, which means at most a fraction of $r$ nodes are able to collude with each other.
Theorem 2.1
Under Assumption 1, Assumption 2 and Assumption 3, the asserter has a dominant strategy to act honestly if and only if
$R_{A}+pS-(1-p)U_{1}-C+(1-r^{n}-(1-r)^{n})pC>pr^{n}(U_{2}+S-R_{V}).$
Proof
The proof is provided in the Appendix 0.A.
Corollary 1
If
$R_{A}+pS-(1-p)U_{1}-C>pr^{n}(U_{2}+S-R_{V}),$
the asserter has a dominant strategy to act honestly.
Proof
We can prove this Corollary by the inequality of $1-r^{n}-(1-r)^{n}\geq 0$ and Theorem 2.1.
3 Application to Decentralized AI Inference Network Design
With the rapid growth in AI, centralized AI servers and corporations are beginning to encounter significant limitations. For instance, it is nearly impractical for a limited number of centralized servers to fulfill the burgeoning global demand for AI inference. Consequently, the concept of a decentralized AI inference network has gained popularity in recent times. Such a network leverages the computational power of a wide array of individual providers who contribute to the AI server pool in a permissionless manner. A well-designed decentralized AI inference network is able to balance the supply with the demand for AI inference capabilities.
In a decentralized network, simply executing a model and trusting its output is insufficient. Consider a scenario where the network is requested to analyze a complex governance issue using a sophisticated model, such as Llama2–70B. There arises a critical question: how can one be certain that the network is not employing a less capable model, like Llama2–13B, thereby delivering subpar analysis while surreptitiously profiting from the discrepancy?
In a centralized setting, one might place their trust in established companies like OpenAI, largely because such entities have a reputation to uphold. Moreover, the quality of a Language Learning Model tends to be self-evident to some degree. However, in a decentralized environment, trust cannot be taken for granted—it must be established through rigorous verification.
SpML is ingeniously designed to excel in the verification process, thereby ensuring the integrity of the system with a negligible increase in computational overhead for security purposes. It stands out as a particularly rational solution when compared to other methods that have been proposed. The following sections will detail the application of PoSP protocol in establishing a robust decentralized AI inference network.
3.1 System Architecture
Deterministic ML Execution. To combat the inherent inconsistencies caused by floating-point calculations in ML, our network implements fixed-point arithmetic and software-based floating-point libraries. To fix the randomness, both the asserter and the validator will be assigned the same random seed. This approach ensures uniform, deterministic ML executions, enabling the use of a deterministic state transition function for the ML process, enhancing reliability in decentralized environments.
Stateless Design in ML Inference. While applications such as inference might appear stateful to users—due to the ability to engage in ongoing interactions—they are fundamentally stateless. In our framework, each query is treated as independent; any necessary historical context is encapsulated within each new request, thereby maintaining statelessness throughout the ML process.
Permissionless Network Participation. In our network, anyone can join the network, gain full access to the ML model $f(x)$, and contribute by running an AI server. The function $f(x)$, pivotal for the AI inference process, is established as common knowledge within the network. This inclusivity ensures that the model is capable for validation, promoting the security of the network.
Off-chain Operations. Given the extensive usage inherent in AI inference networks, it is impractical to process or record every AI inference outcome on the blockchain due to scalability constraints. Instead, AI inferences are computed off-chain by decentralized servers, which then relay the results along with their digital signatures directly to users, bypassing the on-chain mechanism. This approach significantly reduces the load on the blockchain while ensuring that users receive authenticated and accurate inference results.
On-chain Operations. Critical functions, such as the posting of overall balance calculations after set periods, are conducted on-chain to ensure transparency and security. Additionally, the protocol allows for on-chain handling of challenge mechanism, enabling disputes or anomalies detected off-chain to be resolved transparently and securely within the blockchain framework. This ensures both integrity and accountability in the system’s operations.
3.2 SpML Protocol
The spML protocol is designed as follows.
  1. 1.
    The user sends the input to a randomly selected server A with his signature. The base fee is $u$, which is determined by the protocol, and will be burnt.
  2. 2.
    Server A sends the output and its hash back to the user with his signature.
  3. 3.
    With a predetermined probability $p$, the protocol randomly select another server B. Otherwise, A gets the reward $R$, and this round of AI inference concludes.
  4. 4.
    If selected, user sends the same input to server B with his signature.
  5. 5.
    Server B sends the output and its hash back to the user with his signature.
  6. 6.
    The user compares the hash of A’s and B’s output. If they are the same, user accepts the output and gets a pre-determined discount $d<1$ of his base fee, which equals to $d\cdot u$, back to his account, both of A and B gets the reward $R$, and this round of AI inference concludes.
  7. 7.
    Otherwise, the user broadcast A’s and B’s hash to the network. Server A and B also broadcast the user’s input to the network, to prevent the user sends different input to them.
  8. 8.
    All the nodes vote for the behavior of user, sever A and server B. A result will be considered as valid if at least $2/3$ of the nodes vote on it. If the user is dishonest, the user will be slashed $S_{u}$. Otherwise, the user will get the transaction fee and the compensation $u+S/3$, the dishonest server will get slashed $S$ and the honest server will get rewarded $R$.
The user and the voting procedure effectively acts as the orchestrator and arbitration procedure respectively, as in PoSP in Section 2.1, due to the comprehensive participation of network nodes. The integrity of the voting outcome is upheld with the same level of trust as that observed in Ethereum Layer 1. Unlike conventional approaches that may require all nodes to recompute the function $f(x)$ during fraud detection processes, spML protocol simplifies this by necessitating only hash comparisons in most cases. This task parallels the execution of a rudimentary smart contract, incurring little computational overhead. Consequently, our network does not experience significant additional load during challenge mechanism operations.
3.3 Analysis
Property 2
The rational user will never collude with the server in the network, and the dominant strategy for the user is always act honestly.
This is because it is always impossible to slash the honest node. Then we can assume that the user will not collude with any server in the network. Moreover, if the challenge mechanism is triggered, the rational user will indeed send his input to another server B, because he has the utility to gain the transaction fee discount and get to know the true outcome of his request.
Given Property 2, we can assume that the user will always act honestly. In this case, the orchestrator, i.e. the user, is honest, and the spML protocol in Section 3.2 will reduce to the PoSP protocol in Section 2.1.
Proposition 1
If
$p>\frac{C}{(1-r)(R+S)},$
the system has a unique Nash Equilibrium in pure strategies, where every participant perform honestly.
Proof
After plugging $n=1$, $R_{A}=R_{V}=R$, $U_{1}=R$ and $U_{2}=2R$ in Theorem 2.1, we can get this result.
As you can see from Proposition 1, the numerator equals to the computational cost for running one ML model, which is considered to be much less than the denominator. This means if we design the value of the reward and penalty appropriately, we only need little extra computational overhead to guarantee the security of the network.
3.4 SpML vs. Existing Decentralized AI Solutions
In this section, we compare spML with the two prevalent methodologies in decentralized AI networks: optimistic fraud proof based approach (opML) and zero knowledge proof based approach (zkML).
opML. Contrary to the heavy cryptographic reliance of zkML, opML adopts a fundamentally different strategy based on dispute resolution mechanisms. The optimistic approach presupposes that participants will act honestly, given the economic disincentives for fraudulent behavior. In the rare event of disputes, opML provides mechanisms for challenge and resolving fraudulent claims, ideally without necessitating heavy computational verification for every transaction. Nevertheless, the reliance on economic incentives and dispute resolution may introduce vulnerabilities for network security.
zkML. At its core, zkML leverages zero-knowledge proofs. In the context of decentralized AI, zkML ensures that computations can be verified for correctness without revealing the underlying data or the specifics of the computation. This characteristic is particularly advantageous for applications requiring stringent data privacy measures. However, the sophistication and computational intensity of generating zero-knowledge proofs present challenges in terms of scalability and accessibility.
Aspect opML zkML spML
Security More vulnerable to fraudulent activity High security through cryptographic proofs High security through economic incentives
Delays Potential delays in dispute resolution Delays due to proof generation Almost no delay due to asserter will act honestly
Scalability Highly scalable Limited by computational overhead of proof generation Highly scalable
Simplicity Simple, unless for fraud proof Complex due to zk proofs Simple
Overhead Low computational overhead, unless in the case of disputes High computational overhead due to the nature of cryptographic proof generation Low computational overhead, unless in the case of disputes which never happens if everyone is rational
The table above compares opML, zkML and spML as follows.
Security. OpML is recognized as being more vulnerable to fraudulent activity, a vulnerability that can arise from insufficient validation of results. Typically, users themselves may initiate a fraud proof if the results they receive appear aberrant. Nevertheless, it presents a challenge for users to discern the accuracy of these results, particularly when a server might employ a similar yet less sophisticated model to produce the output. Conversely, zkML boasts robust security due to its use of cryptographic proofs. SpML can also achieve high-security just like zkML, but achieves this through economic incentives. In spML, the initiation of challenge mechanism is an automated process managed by the protocol itself, rather than relying on the assumption that there will be at least one external validator, as is the case with opML. This intrinsic mechanism enhances the overall security of our system.
Delays. In opML, delays exist due to the challenge period, during which a transaction can be challenged with a fault proof. This is a drawback in scenarios requiring real-time results. zkML faces inherent significant delays due to the computational overhead in proof generation. SpML is designed to mitigate delay issues altogether. Even if the challenge mechanism is triggered, the user does not need to wait for the challenge procedure, because the result from the asserter is trustable due to the unique Nash Equilibrium in pure strategies.
Scalability. OpML is recognized for its scalability, especially when disputes are minimal, suggesting a lightweight protocol suitable for large-scale applications. ZkML’s scalability is hampered by the heavy computational load required for proof generation. In contrast, spML is presented as highly scalable as well, which can handle extensive network activity without significant degradation in performance.
Simplicity. The complexity of ZK proofs contributes to the high complexity of zkML, possibly making it less accessible for broader implementation. For opML, it is usually considered to be simple. However, its implementation can become complex when applied to fraud proof scenarios. In contrast, spML maintains consistent simplicity in implementation due to its arbitration procedure based on voting. This simplicity enhances ease of integration and maintenance, potentially facilitating widespread adoption.
Overhead. OpML claims low computational overhead, with the caveat that opML may incur higher overhead during disputes. ZkML’s approach results in high computational overhead due to cryptographic processes. SpML also has a low computational overhead. During the challenge mechanism, spML still has a low computational overhead. This is because in spML, the validators only need to compare some hashes in most cases. Only when the hashes do not match during the challenge mechanism, spML may incur high overhead during arbitration, but the arbitration never happens if every node is rational.
3.4.1 Empirical Evaluation
For this part, we use empirical evaluation to further compare opML, zkML and spML.
For zkML, existing solutions, as demonstrated in [4, 8], indicate that generating a proof for a nanoGPT model with $1$M parameters takes approximately $16$ minutes. However, for more advanced models like Llama2-70B, which possesses $70,000$ times more parameters than nanoGPT, it is reasonable to expect that generating a single proof could take several days or weeks. Consequently, employing zkML in a decentralized AI inference network may not be practical given the extended time requirements.
In the opML scenario, when the validator initiates the fraud proof procedure and detects the fraud, we assume the penalty for the malicious server is $S$, and the net gain for the validator is $R_{C}$, accounting for the difference between the reward and the cost of initiating the fraud proof procedure.
Strategy Server Fraud Server Not Fraud
Validator Check $R_{C},-S$ $-C,R-C$
Validator Not Check $0,R$ $0,R-C$
The table above gives a game in a bimatrix format, where the first number in each pair represents the utility to the validator, and the second number represents the utility to the server. We can calculate the probability for the undetected fraud is $(S+R-C)C/[(S+R)(R_{C}+C)]$ by the mixed strategy Nash Equilibrium, similar to the approach detailed in [14]. Assuming $R_{C}=100C$, $R=1.2C$ and $S=150C$, the calculated probability of undetected fraud is $0.98\%$. This implies that if you request AI inference $50$ times per day, you can expect, on average, one undetected fraud approximately every $2$ days.
In contrast, in the spML scenario, assuming the asserter controls a fraction $r=10\%$ of the network, by Proposition 1, the probability of triggering the challenge mechanism is $0.735\%$. This translates to only $0.735\%$ additional computational overhead in spML, enabling us to completely avoid fraud and eliminate the need for fraud proof procedures, if all nodes are rational. Hence, spML is clearly the superior choice.
4 Conclusions and Future Extensions
In this study, we have introduced a robust PoSP protocol applicable across all kinds of scenarios, such as decentralized AI inference network, distinguished by its exceptional security features. Central to its design is the implementation of a unique Nash Equilibrium in pure strategies, ensuring that all rational participants within the network adhere to honest behavior. This protocol demonstrates superior performance when compared to existing methodologies across various applications.
Looking ahead, further exploration into the application of the PoSP protocol within Layer 2 architectures holds promise, particularly by employing our method of sampling multiple nodes to recompute results. This approach can lead to a unique Nash Equilibrium in pure strategies, where every participant acts honestly, directly addressing and potentially solving the concerns highlighted in [14]. Following this, there is significant potential for utilizing the PoSP protocol in the development of verification mechanisms within Actively Validated Services (AVS), enhancing security and reliability in EigenLayer. This exploration could lead to innovative applications that leverage the strengths of PoSP in ensuring robust, scalable, and secure systems.
Appendix 0.A Proof of Theorem 2.1
First, we consider the expected payoff of the asserter if he does not commit fraud. If the asserter does not commit fraud, the expected payoff is
$(1-p)(R_{A}-C)+p\sum_{i=0}^{n}\binom{n}{i}r^{i}(1-r)^{n-i}\left(i\frac{R_{V}}{n}+R_{A}-C\right).$
This is because when challenge mechanism is triggered, if $i$ out of $n$ selected validators collude with the asserter,the collusions can have the payoff $iR_{v}/n+R_{A}-C$, because the colluders are able to obtain the outcome of $f(x)$ directly from the asserter instead of recalculating it themselves.
Then, we consider the expected payoff of the asserter if he commits fraud. If the challenge mechanism is not triggered, the asserter is capable of compromising the integrity of the system and get a payoff of $U_{1}$. However, if the challenge mechanism is triggered, by Property 1, only when all of the $n$ selected validators collude with the asserter, the fraud will not be detected, and the asserter can earn $U_{2}$. If $1\leq i<n$ out of $n$ selected validators collude with the asserter, their optimal strategy is to act honestly and report the fraud to get the reward for validation rather than be slashed. Thus, the expected payoff is
$(1-p)U_{1}+pr^{n}U_{2}+p\sum_{i=0}^{n-1}\binom{n}{i}r^{i}(1-r)^{n-i}\left(i\frac{R_{V}}{n}-C-S\right).$
Hence, the system having a unique Nash Equilibrium in pure strategies is equivalent to the scenario where the asserter can obtain a greater profit if he does not commit fraud, which is
$(1-p)(R_{A}-c)+p\sum_{i=0}^{n}\binom{n}{i}r^{i}(1-r)^{n-i}\left(i\frac{R_{V}}{n}+R_{A}-C\right)$
$>$ $(1-p)U_{1}+pr^{n}U_{2}+p\sum_{i=0}^{n-1}\binom{n}{i}r^{i}(1-r)^{n-i}\left(i\frac{R_{V}}{n}-C-S\right).$
By rearranging this inequality, we can get
$R_{A}+pS-(1-p)U_{1}-C+(1-r^{n}-(1-r)^{n})pC>pr^{n}(U_{2}+S-R_{V}),$
which coincides Theorem 2.1.